# A decoder-only foundation model for time-series forecasting

### v1

@@ Abstract
Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Our model is based on pretraining a patched-decoder style attention model on a large timeseries corpus, and can work well across different forecasting history lengths, prediction lengths and temporal granularities. Author names are listed in alphabetical order.
@@ 1 Introduction
In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets with different temporal granularities, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets). Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time. The key elements of our foundation model are twofold: 1) a time series corpus built mostly using Google Trends [^1] https://trends.google.com, that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus. Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (225M parameters) and pretraining data size (1B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data.
@@ 2 Related Work
[26] show how to use the the GPT-2 backbone [16] for various tasks including time-series forecasting. [4] is a follow up works along the same lines. Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset. For instance Table-18 [26] shows M4 to M3 transfer. The rest of the two papers are mostly focused on fine-tuning and to the best of our knowledge they do not train a single foundation model that shows out of the box zero-shot performance on a variety of datasets. To the best of our knowledge, the very recent work in TimeGPT-1 [6] is the only known parallel work on foundation model for time-series. However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
@@ 3 Problem Definition
The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points. Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy like notation for indices. Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$. Note that since we are building a one-fits-all model we cannot have dataset specific dynamic or static covariates during training time. However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$. See Appendix A.1 for details. Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$. The task is then to learn a capable foundation model that can map any time-series context to horizon, given by
@@
$\mathrm{MSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)=\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}.$ (2)
@@ 4 Model Architecture
A foundation model for time-series forecasting should be able to adapt to variable context and horizon lengths, while having enough capacity to encode all patterns from a large pretraining datasets. Transformers have been shown to be able to adapt to different context lengths in NLP [16]. Inspired by the success of patch based modeling in the recent long horizon forecasting work [13] we also chose to breakdown the time-series into patches during training. However, there are several key differences in our foundation model architecture, the primary one being that our model is trained in decoder-only mode [9]. We will now describe the key parts of our architecture and training methodology illustrated in Figure 1.
[…]
[width=]fts_new_v2.pdf
@@
Input Layers. The job of the input layers is to preprocess the time-series into input tokens to the transformer layers. We first break the input into contiguous non-overlapping patches. Then each patch (along with optional date derived features for that patch) is processed by a Residual Block into a vector of size model_dim. The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [5].
In other words, the inputs $\mathbf{y}_{1:L},{\bm{x}}_{1:L}$ are broken down into patches of size input_patch_len ($p$). The $j$-th patch can be denoted as $\tilde{\mathbf{y}}_{j}=\mathbf{y}_{p(j-1)+1:pj}$ and $\tilde{\mathbf{x}}_{j}=\mathbf{x}_{p(j-1)+1:pj}$. Then the $j$-th input token to the subsequent transformer layers can be denoted as,
$\mathbf{t}_{j}=\texttt{InputResidualBlock}(\tilde{\mathbf{y}}_{j},\tilde{\mathbf{x}}_{j})+\texttt{PE}_{j}$ (3)
@@
Stacked Transformer. The bulk of the parameters in our model are in $n_{l}$ transformer layers stacked on top of each other. Each of these layers have the standard multi-head self-attention (MHA) followed by a feed-forward network (FFN). The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads). We set the hidden size of the FFN’s to be equal to model_dim as well. We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token). This can be described by the equation
$\mathbf{o}_{j}=\texttt{StackedTransformer}\mathopen{}\mathclose{{}\left(\mathbf{t}_{1},\cdots,\mathbf{t}_{j}}\right),~{}~{}~{}\forall j\in[N].$ (4)
@@
Inference. The trained network can be used to produce forecasts for any horizon using auto-regressive decoding similar to large language models. Given an input $\mathbf{y}_{1:L}$ (assume $L$ is a multiple of $p$ for simplicity) it can first predict $\hat{\mathbf{y}}_{L+1:L+h}$. Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on.
We name our model Pretrained Decoder for Time-series (PreDcT).
@@ 5 Empirical Results
We evaluate our model in zero-shot settings on well known public datasets against state-of-the-art supervised forecasting baselines. We show that a single pretrained model can come close or surpass the performance of baselines on the benchmarks even when the baselines are specially trained or tuned for each specific task. Subsequently, we perform ablation studies that justify different choices made in our foundation model architecture.
@@ 5.1 Zero-shot Evaluation
We are interested evaluating the performance of a single foundation model pretrained on a large time-series dataset on different target datasets never seen before by the model or, in other words, zero-shot evaluation. To test the generalization performance of our baselines, we choose target datasets covering a variety of data domains, scales and time granularities. We will first discuss the pretraining and target datasets, along with the baselines we use. Then we demonstrate the (Z)ero-(S)hot performance of our model PreDcT(ZS) compared against state-of-the-art models trained directly on the target datasets.
Pretraining Data. We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model. It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model. In this paper, we find that Google Trends [^2] https://trends.google.com can provide a time series corpus that is ideally suited for pre-training our foundation model. Google Trends captures search interest over time for millions of queries. We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022. We use the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset. The date ranges are Jan. 2018 to Dec 2019 for hourly and Jan. 2007 to Dec. 2021 for the other granularities. Along with the trends data, we also add time series from several other publicly available datasets to our pretraining corpus. We add in all the granularities of the M4 dataset [12] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [28]). M4 has a good mix of granularities with around 100k time-series in total. Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points. In addition, we add all the 15 min granularity traffic time series from [22].
We train on a mixture distribution over these datasets that aims to give sufficient weightage to all granularities. We train with a maximum context length of 512 whenever the length of the time-series allows that. For weekly granularity we do not have sufficiently long time-series therefore a max. context length of 256 is used. For the same reason, max. context length of 64 is used while training on monthly and higher granularity data.
@@
Target Datasets. To benchmark our model’s performance, we choose commonly used forecasting datasets of varying sizes that cover various domains, granularities, context lengths and horizon lengths, to test the generalization power of our foundation model against other baselines. The details are summarized in Table 1.
Dataset # Time-Series Time Length Granularity Context Horizon
ETTm1 7 69680 15 Min. 512 96
ETTm2 7 69680 15 Min. 512 96
ETTh1 7 17420 1 Hour 512 96
ETTh2 7 17420 1 Hour 512 96
Wiki 115084 803 1 Day 256 56
ILI 7 966 1 Week 96 24
TourismL 555 228 1 Month 64 12
Table 1: Statistics and task definitions of the target datasets.
(Sub)Hourly.  For 15 min. granularity we use the ETTm1, ETTm2 datasets and test all models on the task of predicting a horizon of 96 time-points after seeing a context of size 512. For hourly granularity, we choose the ETTh1, ETTh2 datasets and test all models on the same task as the 15 min. datasets. The datasets and the context, horizon pair have been widely used in long-term forecasting benchmarks [28, 13]. Note that we used the more challenging original, unscaled versions of these datasets in order to test our model’s zero-shot performance on time-series of different scales.
Daily.  For the daily granularity, we use the Wikipedia web-traffic dataset from the corresponding Kaggle competition  [^3] https://www.kaggle.com/code/muonneutrino/wikipedia-traffic-data-exploration. It has 115k time-series with over two years of data if we exclude the time-series with missing values. The dataset contains web-traffic to Wikipedia articles and the task is to predict the web-traffic (in $\log$ scale) on future dates. We choose a context length of 256 to predict a horizon of 56 days (8 weeks). This dataset has been used in prior multivariate forecasting papers such as like [18].
Weekly.  We use the ILI dataset [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html that collects the number of patients and influenza-like illness ratio in a weekly frequency. We use a context length of 96 to predict 24 weeks into the future. This is one of the configurations used in long-term forecasting papers like [28].
Monthly.  We choose TourismL (Tourism Large) [21] as one of the target datasets. It contains monthly tourist visit data in Australia that has been grouped into various regions. It consists of 555 time-series with very different scales. We choose the task of predicting a 12 month horizon given a context length of 64.
All the target datasets are divided into train:validation:test splits (periods) chronologically with the proportions being 7:1:2. We evaluate the models based on metrics resulting from rolling windows in the test period. Specifically, PreDcT(ZS) solely predicts in the test period as it has already been pretrained. The supervised learning models (per dataset) are trained on the train part with the hyper-parameters being tuned using the validation split. Then they predict in the test period for a head-to-head comparison with the zero-shot PreDcT(ZS).
Baselines. We compare our model against three recently published state-of-the-art supervised forecasting models PatchTST [13], TiDE [5] and FEDFormer [25] as well as the popular N-BEATS model [14]. Note that these models have already been shown [25, 28] to be superior to common statistical forecasting methods such as Prophet [19] and ARIMA; hence we do not include them in our baselines. We train the above supervised models on the train split of each target dataset and measure their performance on the corresponding test split. For these supervised models we report the best metrics among models trained with and without date-derived features. See Appendix A.2 for the hyper-parameters used for each dataset. We compare our zero-shot metrics from PreDcT(ZS) to these state-of-the-art supervised metrics, which we denote by PatchTST(S), TiDE(S), N-BEATS(S) and FEDFormer(S) in our results below.
In PreDcT(ZS), we set input_patch_len=32, output_patch_len=128. We train a model with about 225M parameters that uses 16 head multi-head attention in each transformer layer.
Results. In Table 2 we present the main results on all our target datasets. We report normalized metrics NRMSE and WAPE, that are proportional to MSE and MAE, and are defined (for each time series) as
$\mathrm{NRMSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\sqrt{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}},$

### v2

@@ Abstract
Author names are listed in alphabetical order.1. Google Research, Correspond to senrajat@google.com
Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Our model is based on pretraining a patched-decoder style attention model on a large time-series corpus comprising both real-world and synthetic datasets. Experiments on a diverse set of previously unseen forecasting datasets suggests that the model can yield accurate zero-shot forecasts across different domains, forecasting horizons and temporal granularities.
Machine Learning, ICML
[…]
@@ 1 Introduction
In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets). Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time. The key elements of our foundation model are twofold: 1) a large-scale time series corpus built using both synthetic and real-world (mostly using time series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/), that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus.
Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (100B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data. Our work also suggests that unlike recent work [9] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a fraction of its costs.
@@ 2 Related Work
There has been some recent work on re-using or fine-tuning large language models for time-series forecasting. In particular,  [9] benchmarks pretrained LLMs like GPT-3 and LLaMA-2 for zero-shot forecasting performance. As we show later, our model obtains much superior zero-shot performance at a tiny fraction of these model sizes. Zhou et al. [36] and [5] show how to fine-tune a GPT-2 [24] backbone model for time-series forecasting tasks. Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset. With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on a target dataset, and not on pretraining a single foundation model that shows out-of-the box zero-shot performance on a variety of datasets. To the best of our knowledge, the very recent work in TimeGPT-1 [7] is the only other parallel work on a zero-shot foundation model for time-series forecasting. However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
@@ 3 Problem Definition
The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points. Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy-like notation for indices. Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$. Note that since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time. However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$. Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$. The task is then to learn a capable foundation model that can map any time-series context to horizon, given by
@@
$\mathrm{MSE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)$
$=\frac{1}{H}\mathopen{}\mathclose{{\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}}\right\|_{2}^{2}.$ (2)
@@ 4 Model Architecture
A foundation model for time-series forecasting should be able to adapt to variable context and horizon lengths, while having enough capacity to encode all patterns from a large pretraining datasets. Transformers have been shown to be able to adapt to different context lengths in NLP [24]. However, there are several time-series specific design choices. The main guiding principles for our architecture are the following:
Patching. Inspired by the success of patch based modeling in the recent long horizon forecasting work [19] we also choose to breakdown the time-series into patches during training. A patch of a time-series is a natural analogues for a token in language models and have been shown to improve performance. Moreover this improves inference speed as the number of tokens being fed into the transformer is reduced by a factor of the patch length. On the other hand, increasing the patch length all the way to the context length moves us away from decoder-only training and the efficiencies that come with it. We delve into this further in Section 6.2.
Decoder-only model. A key difference between our architecture and PatchTST [19] is that our model is trained in decoder-only mode [15]. In other words, given a sequence of input patches the model is optimized to predict the next patch as a function of all past patches. Similar to LLM’s this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
Longer output patches. In LLM’s the output is always generated in an auto-regressive fashion one token at a time. However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy that multi-step auto-regressive decoding [33]. But the latter is not possible when the horizon length is not known a prior like in the case of zero-shot forecasting, which is our primary goal.
We propose a middle ground by allowing the output patches for prediction to be be longer than the input patches. As an example, suppose the input patch length is 32 and output patch length is 128. During training, the model is simultaneously trained to use the first 32 time-points to forecast the next 128 time-steps, the first 64 time-points to forecast time-steps 65 to 192, the first 96 time-points to forecast time-steps 97 to 224 and so on. During inference, suppose the model is given a new time-series of length 256 and tasked with forecasting the next 256 time-steps into the future. The model will first generate the future predictions for time-steps 257 to 384, then condition on the initial 256 length input plus the generated output to generate time-steps 385 to 512. On the other hand, if in our model the output patch length was equal to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above.
Patch Masking. If we use patches naively, the model might only learn to predict well for context lengths that are multiples of the input patch length. Therefore we make a careful use of masking during training. Parts of patches as well as entire patches from the beginning of the context window can be masked in a data batch. We employ a specific random masking strategy during training that helps to model see any all possible context lengths starting from 1 to a maximum context length.
Now that we have mentioned the guiding principles, we next formally describe each component of our architecture that is illustrated in Figure 1.
@@
Input Layers. The job of the input layers is to preprocess the time-series into input tokens to the transformer layers. We first break the input into contiguous non-overlapping patches. Then each patch is processed by a Residual Block into a vector of size model_dim. Along with the input, we also supply a binary padding mask $\mathbf{m}_{1:L}$ where 111 denotes that the corresponding input in $\mathbf{y}_{1:L}$ should be ignored and vice-versa. The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [6].
In other words, the inputs $\mathbf{y}_{1:L}$ are broken down into patches of size input_patch_len ($p$). The $j$-th patch can be denoted as $\tilde{\mathbf{y}}_{j}=\mathbf{y}_{p(j-1)+1:pj}$. Similarly the mask can also be patched as $\tilde{\mathbf{m}}_{j}=\mathbf{m}_{p(j-1)+1:pj}$. Then the $j$-th input token to the subsequent transformer layers can be denoted as,
$\mathbf{t}_{j}=\texttt{InputResidualBlock}(\tilde{\mathbf{y}}_{j}\odot(1-\tilde{\mathbf{m}}_{j}))+\texttt{PE}_{j}$ (3)
@@
Stacked Transformer. The bulk of the parameters in our model are in num_layers ($n_{l}$) transformer layers stacked on top of each other. Each of these layers have the standard multi-head self-attention (SA) followed by a feed-forward network (FFN). The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads). We set the hidden size of the FFN’s to be equal to model_dim as well. We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token). This can be described by the equation
$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$
$\forall j\in[N].$ (4)
In Equation (4), $\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked. All patches that are masked out completely are not attended to by the causal self attention.
@@
Training. We train the model with standard mini-batch gradient descent in decoder only fashion, that goes through all windows for a time-series and across time-series. The only non-standard part is the way we sample the mask during training. For each time-series in the batch, we sample a random number $r$ between 0 and $p-1$. Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask our a fraction of the first input patch. However, this is sufficient to cover all input context lengths from 1 to the maximum training context length.
Let us say that the maximum context length is 512 and $p=32$. Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized predict after seeing 28=32428=32-428 = 32 - 4 time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing 28+3228+3228 + 32 time-points, and so on. When this argument is repeated for all such $r$’s, we can see that the model has seen all possible context lengths from 111 to 512512512.
@@
Inference. The trained network can be used to produce forecasts for any horizon using auto-regressive decoding similar to large language models. Given an input $\mathbf{y}_{1:L}$ (assume $L$ is a multiple of $p$ for simplicity) it can first predict $\hat{\mathbf{y}}_{L+1:L+h}$. Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on. If $L$ is not a multiple of $p$ we can append some zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.
We name our Time-series Foundation Model as TimesFM.
5 Pretraining Details
We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model. It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model. The bulk of real data used to train our models come from Google trends  [^3] https://trends.google.com and Wiki Pageview statistics  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics. In summary the main data sources are:
Google Trends. Google Trends captures search interest over time for millions of queries. We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022. We use the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset. The date ranges are Jan. 2018 to Dec. 2019 for hourly and Jan. 2007 to Dec. 2021 for the other granularities.
Wiki Pageviews. Wiki Pageviews captures the hourly views of all Wikimedia pages. We download all pageview data from Jan. 2012 to Nov. 2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out the resulting time-series with excessive zeros. The final corpus contains O(100B) time points.
Other real data sources. Along with the wiki and trends data, we also add time-series from several other publicly available datasets to our pretraining corpus. We add in all the granularities of the M4 dataset [17] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [34]). We also add the 10 minute granularity Weather dataset used for evaluations in [34]. M4 has a good mix of granularities with around 100k time-series in total. Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points. In addition, we add all the 15 min granularity traffic time series from [29].
Synthetic. Another major component of our pretraining data is of synthetic origin. We create generators for ARMA [18] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions. A synthetic time-series can be an additive combination of 1 or more of these processes. We create 3M synthetic time-series each of length 2048 time-points. More details are presented in Appendix A.4.
Dataset Mixing. We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities. The training mixture is 50% real data and 50% synthetic, with the real data mixture providing equal weights to all hourly + sub-hourly, daily, weekly, and monthly datasets. We train with a maximum context length of 512 whenever the length of the time-series allows that. For weekly granularity we do not have sufficiently long time-series therefore a maximum context length of 256 is used. For the same reason, a maximum context length of 64 is used while training on $\geq$ monthly granularity data.
(a) Monash Archive [8]
(b) Darts [10]
(c) ETT (Horizons 96 and 192) [34]
Figure 2: We report average performance in three groups of datasets. In all figures, the lower the metric the better and the error bars represent one standard error. Note that among the baselines onlye TimesFM and llmtime are zero-shot. In (a) we report results on the Monash datasets. Since the datasets have different scales, we average the MAE’s scaled by the MAE of a naive baseline. We can see that TimesFM is among the top-3 models. In (b), we report the similarly scaled MAE on the Darts benchmarks. TimesFM is within significance of the top of the best performing method which is ARIMA in this case. Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones. Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total. TimesFM and PatchTST are the best performing models in this case.
@@ 6 Empirical Results
We evaluate our model in zero-shot settings on three groups of well known public datasets against the best performing baselines for each group. These datasets have been intentionally held out from our pretraining data. We show that a single pretrained model can come close or surpass the performance of baselines models on the benchmarks even when the baselines are specially trained or tuned for each specific task. Subsequently, we perform ablation studies that justify different choices made in our architecture.
@@ 6.1 Zero-shot Evaluation
To benchmark our model’s performance, we choose three groups of commonly used forecasting datasets that cover various domains, sizes, granularities, and horizon lengths: Darts [10], Monash [8] and Informer datasets [34], to test the generalization power of our foundation model against other baselines.
In all cases, we report performance on the official metrics and scalings of the datasets. We present a summary of the results below - more detailed experimental results can be found in Appendix A.2.
Monash [8]. Monash archive is a collection of 30 datasets of different training and prediction lengths that covers granularities ranging from minutes to years and domains including finance, demand forecasting, weather and traffic. The archive reports four official metrics for several statistical baselines such as Exponential Smoothing(ETS) and ARIMA, as well as supervised ML baselines like CatBoost [23], DeepAR [25] and WaveNet [20]. Following llmtime [9] we start from the Monash Huggingface repository  [^5] https://huggingface.co/datasets/monash_tsf and filter out the datasets that contain missing values. This leaves us with 18 datasets.
Out of the four official metrics, following prior work [9], we report our performance in terms of mean MAE (see Appendix A.1). As the datasets have massively different scales, for each dataset we normalize the metric by the metric achieved by a naive baseline that just constantly predicts the last value in the context for each time-series. Then the scaled MAE’s are averaged across all datasets. The scaled aggregation was also used in [9].
The mean scaled MAE across all datasets is plotted in Figure 2(a) along with standard error bars. We compare the performance of TimesFMwith the baseline models implemented in Monash, and the zero-shot llmtime [9] model that uses GPT-3 [24] with a specific prompting technique. Note that the zero-shot models are marked as (Zero-Shot). TimesFM is among the top 3 models even though we never trained on these datasets. It outperforms some deep supervised models like DeepAR and FFNN, and improves on llmtime’s performance by more than 10%.
Darts [10]. This is a collection of 8 univariate datasets which include interesting seasonalities and additive+multiplicative trends. We report performance of several baselines implemented in the Darts package like TCN [14], N-HiTS [4] and N-BEATS [21]. All these baselines are supervised. As before, we also report zero-shot forecasting results from llmtime [9] using GPT-3 [24]. Other supervised baselines implemented in [9] like SM-GP [31] and seasonal ARIMA [18] are also added.
We report the official metric for this dataset group that is MAE for each individual dataset in Appendix A.2. In Figure 2(b), we present the average scaled MAE across all 8 datasets, as we did for the Monash datasets. TimesFM is within statistical significance of the best model that is seasonal ARIMA in this case. Note that since there are only 8 individual time-series in this dataset group, the standard errors are not sharp and therefore does not provide a clear ordering among the models.
Informer [34]. The Informer datasets have been widely used for benchmarking various supervised long-horizon forecasting methods. A few of these datasets are used in pretraining, so we focus on the other datasets in this collection (ETTm1, ETTm2. ETTh1 and ETTh2) related to electricity transformer temperatures over a two year period in 111 hour and 151515 minutes granularities. Note that the long horizon baselines usually report rolling validation results on the test set which would amount to millions of tokens for evaluating llmtime [9] and would be too expensive. Therefore, following the latter paper, we compare all methods on the last test window. Also, it is reasonable to directly average the MAE for these datasets since the results are reported on standard normalized dataset (using the statistics of the training portion).
We consider the task of predicting horizon length 96 and 192, given a context length of 512 for all methods. The MAE averaged over all 8 tasks (4 datasets with two horizons each) is presented in Figure 2(b). TimesFM performs the best and the supervised PatchTST [19] baseline (which is a state-of-the-art long horizon deep forecasting method) is within significance of it. The other long horizon methods are quite a bit worse even though they have been trained these datasets. llmtime(ZS) is better than FEDFormer but worse than PatchTST in a statistically significant way.
6.2 Ablation
Next, we perform several ablation studies that inform the design decisions we made for our model architecture.
Scaling. Performance curves with respect to number of parameters in a model have been a keenly studied area in the context of LLMs.  Kaplan et al. [12] established a power law like relationship between the number of parameters in a language model and its downstream performance i.e the more the number of paramaters the better the performance. However,  Hoffmann et al. [11] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset. We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pretraining dataset and till the similar number of iterations.
We provide the mean scaled MAE results on the Monash datasets in Figure 3. It can be clearly seen that the errors decrease monotonically with the number of parameters, thus pointing to a similar observation as [12], but in the context of time-series.
Figure 3: Average scaled MAE on Monash datasets for three different TimesFM model sizes. We can see that the performance improves with increasing model size.
Autoregressive decoding. In recent long-term forecasting works [33, 19, 6] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks. For a foundation model the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons. However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps. This was one of the key decisions in the design of TimesFM, that is quite different from LLMs. In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets. In Figure 4, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128. The former has to perform 16 autoregressive steps while the latter has to do only 4. It can be clearly seen that having a larger output_patch_len helps in this case.
Figure 4: Ablation with respect to output patch length for the task of predicting 512 steps into the future, in the same setting as that of Section 6.1. We report the average across all ETT 4 datasets.
Input patch length.  The size of input_patch_len represents an important trade-off. We have typically seen that increasing its value from 8 to 32 increases performance but having too high a input_patch_len is impractical since that makes the model shift from decoder only training more towards encoder-decoder style training. Note that in the ”Training” paragraph of Section 4, we describe the mask sampling strategy to support any context length. If in the extreme case $p$ is set the maximum context length we have to individually sample all possible context windows from 1 to maximum context length, which would be required for encoder-decoder style of training.
In Figure 5 we show the mean scaled MAE TimesFM(ZS) - 70M model with input_patch_len=8 on Monash datasets, which is clearly worse than our original model that uses input_patch_len=32.
Figure 5: Average scaled MAE for our 70M models on Monash datasets for two different input patch lengths. We can see that the performance improves with increasing patch size.
6.3 Visualizing Forecast
We conduct a visual inspection of the forecasts generated by TimesFM(ZS) on the benchmark datasets. Our inspection in particular emphasizes time-series exhibiting both trend and seasonal components readily interpretable by humans. As illustrated in Figure 6, TimesFM(ZS) effectively captures these subtle characteristics within both the trend and seasonal patterns of the depicted time-series.
Figure 6: Forecasts visualized on Darts and Monash datasets. Plots on the right zoom in the forecasting parts from the left.
7 Conclusion
In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time series data. This model is trained on real-world and synthetic datasets comprising around 100B timepoints using a patched-decoder style attention architecture with around 200M parameters.
In future work, we plan to delve into a more theoretical understanding of how such a time-series foundation model can obtain good performance for out-of-distribution data, and also investigate the fine-tuning/few-shot performance of this model.
Appendix A Appendix
A.1 Metrics
The metrics that are used for reporting results in this paper are:
  • MAE [8]
    $\mathrm{MAE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)\\ =\frac{1}{H}\mathopen{}\mathclose{{\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}}\right\|_{1}.$ (7)
  • msMAPE [8]